OpenStudy (anonymous):

Solve the folowing matrix for c3,c4 and c1 \[\huge \left[\matrix{n&&c_3&&c_4&&c_1\\-\frac{20pa^3}{3}&&2a&&1&&-2a\\-6pa^3&&3a&&1&&0\\-2pa^2&&1&&0&&-1} \right]\]

OpenStudy (anonymous):

sorry, I have no idea what that even is! I can toss the link around to people that night tho

OpenStudy (anonymous):

Ok if you could just solve this I would appreciate it \[\huge \left[\matrix{-\frac{20pa^3}{3}&&2a&&1&&-2a\\-6pa^3&&3a&&1&&0\\-2pa^2&&1&&0&&-1} \right]\]

OpenStudy (anonymous):

This is what I did \[\huge \left[\matrix{-\frac{20pa^3}{3}&&2a&&1&&-2a\\-6pa^3&&3a&&1&&0\\-2pa^2&&1&&0&&-1} \right]\] Multiplied top row by 1/(2a), second row by 1/(3a) and third row by -1 to get this new matrix \[\huge \left[\matrix{-\frac{10pa^2}{3}&&1&&\frac{1}{2a}&&-1\\-2pa^2&&1&&\frac{1}{3a}&&0\\2pa^2&&-1&&0&&+1} \right]\] I then added the first row to the last row and replaced the first row with the result to obtain this new matrix \[\huge \left[\matrix{-\frac{10pa^2}{3}+2pa^2&&0&&\frac{1}{2a}&&0\\-2pa^2&&1&&\frac{1}{3a}&&0\\2pa^2&&-1&&0&&+1} \right]\] \[\huge \frac{-10pa^2+6pa^2}{3} = \frac{-4pa^2}{3}(2a)=c_4\] Then I get \[\huge c_4 = -\frac{8pa^3}{3}\] WHICH IS WRONG THE RIGHT ANSWER IS \[\huge c_4 = \frac{4}{3}Pa^3\]

OpenStudy (amistre64):

what do you mean by "solve"?

OpenStudy (anonymous):

its a system of equations

OpenStudy (amistre64):

can you show the equations? and whats the top row represent?

OpenStudy (anonymous):

in a second

OpenStudy (anonymous):

P and a are constants

OpenStudy (amistre64):

an augment is usually, as far as ive seen; written with the coeffs on the left and the =stuff to the right

OpenStudy (anonymous):

Same thing

OpenStudy (anonymous):

You get the picture

OpenStudy (amistre64):

i see it ....

OpenStudy (amistre64):

\[\begin{vmatrix}-\frac{20pa^3}{3}&|&2a&1&-2a\\-6pa^3&| &3a&1&0\\-2pa^2&| &1&0&-1\end{vmatrix}\] \[\begin{vmatrix}-2pa^2&| &1&0&-1\\ -\frac{20pa^3}{-2a.3}&|&-1&1/-2a&1\\-6pa^3/-3a&| &-1&1/-3a&0\end{vmatrix}\] \[\begin{vmatrix}-2pa^2&| &1&0&-1\\ \frac{10pa^2}{3}-2pa^2&|&0&-1/2a&0\\2pa^2-2pa^2&| &0&-1/3a&-1\end{vmatrix}\] \[\begin{vmatrix}-2pa^2&| &1&0&-1\\ \frac{-4pa^2.2a}{3}&|&0&1&0\\0&| &0&-1&-3a\end{vmatrix}\] \[\begin{vmatrix}-2pa^2&| &1&0&-1\\ \frac{-8pa^3}{3}&|&0&1&0\\\frac{-8pa^3}{3}&| &0&0&-3a\end{vmatrix}\] \[\begin{vmatrix}-2pa^2&| &1&0&-1\\ \frac{-8pa^3}{3}&|&0&1&0\\\frac{8pa^2}{9}&| &0&0&1\end{vmatrix}\] \[\begin{vmatrix}\frac{-10pa^2}{9}&| &1&0&0\\ \frac{-8pa^3}{3}&|&0&1&0\\\frac{8pa^2}{9}&| &0&0&1\end{vmatrix}\] maybe .....

OpenStudy (amistre64):

you sure you got your eqs correct?

OpenStudy (anonymous):

No I had one wrong equation all along http://www.wolframalpha.com/input/?i=rref%7B%7B2a%2C-2a%2C-1%2C8a%5E3z%2F3%7D%2C%7B1%2C-1%2C0%2C2za%5E2%7D%2C%7B0%2C3a%2C1%2C-6za%5E3%7D%7D I apologize for wasting your time. Really really sorry

OpenStudy (anonymous):

c1 c3 c4 in that order

OpenStudy (amistre64):

yeah, it helps to have the right inputs :)